# Programmable Cognitive Bias in Social Agents

### v1

@@ 9 Limitation
1. Usability: We choose to evaluate CoBRA using technical benchmarks and demonstrations, standard methods for assessing HCI toolkits Ledo et al. (2018). This paper focuses on exploring the technical feasibility and foundational understanding of the problem. We plan deployment studies with sociologists as the next step to assess CoBRA’s usability Landay (2009).
@@ A.1 Cognitive Bias in LLM-As a Feature, Not a Bug
Traditionally, cognitive bias is defined as systematic deviations from rational judgment that lead to suboptimal decision-making outcomes Unknown (2015); Baron (2007). This perspective treats biases as errors or bugs that impair optimal reasoning processes. Recent studies have documented extensive cognitive biases in LLMs, including confirmation bias, anchoring bias, and framing effects Koo et al. (2023); Boonprakong et al. (2025); Sumita et al. (2025). Research demonstrates that LLMs exhibit amplified cognitive biases in moral decision-making contexts Cheung et al. (2025), leading to a systematic assessment and mitigation of these biases Echterhoff et al. (2024).
However, cognitive bias can also be viewed as an adaptive feature that evolved to facilitate social navigation and cooperation Haselton et al. (2015); Cosmides & Tooby (1994). From this evolutionary perspective, biases serve important prosocial functions, enabling sophisticated social intelligence rather than representing flaws Liu et al. (2025b). However, existing research lacks systematic mechanisms to harness these biases as controllable features for enhancing AI social capabilities.
@@ A.2 LLM Behavioral Control Methods
Prompt engineering has served as the foundational approach to behavioral control, leveraging natural language instructions to guide model behavior using structured prompting strategies Rath et al. (2025); Zhuo et al. (2024). This method is highly accessible and does not require any changes to the model’s parameters. However, designing effective prompts has many pitfalls Zamfirescu-Pereira et al. (2023), and fundamental limitations persist: LLMs exhibit extreme sensitivity to prompt variations, with large accuracy differences from minor changes Sclar et al. (2024).
Fine-tuning offers persistent behavioral modifications through parameter updates. Techniques such as Reinforcement Learning from Human Feedback (RLHF) and Direct Preference Optimization (DPO) have become dominant for model alignment Ouyang et al. (2022); Rafailov et al. (2023). Task vectors enable compositional behavioral modifications through arithmetic operations on weight differences Ilharco et al. (2023). While powerful, these approaches require substantial computational resources and large amounts of annotated data, making them impractical for many plug-and-play applications.
Representation engineering provides a method for behavioral control through direct manipulation of a model’s internal representations during inference Zou et al. (2023). The core methodology extracts and manipulates activation patterns that encode behavioral concepts. Activation steering methods compute steering vectors from contrasting behavioral examples Rimsky et al. (2024), which can be applied during inference to modulate behavior without parameter updates Todd et al. (2024). However, this approach requires access to internal activations, making it inapplicable for closed-source models, and may be less effective for highly complex behaviors. Recent applications include personality control Chen et al. (2025), community-specific preference modeling Li et al. (2025), harmlessness control Arditi et al. (2024), and multi-objective alignment Shi et al. (2024). These methods operate with minimal inference overhead while maintaining base model capabilities.
@@ A.3 LLM-based Social Simulation & Agent Specification
Existing work on LLM-based agent social simulations typically instantiates LLMs as autonomous agents by assigning them scenario-specific prompts, roles, and objectives Park et al. (2023); Chen et al. (2024b). Methodologically, these studies typically follow a pipeline consisting of three core stages: 1. Agent Specification, 2. Social Environment Set-up, and 3. Behavior Observation. During Agent Specification, researchers harness the expressiveness and flexibility of natural language to define agent identity, knowledge, and behavioral tendencies, enabling the creation of contextualized personas. In the Social Environment Set-up stage, agents are embedded in experimental scenarios inspired by social science literature Liu et al. (2025b), or in constructed social environments with specific rules and resources Zhou et al. (2024); Park et al. (2023); Chen et al. (2024b), such as sandbox simulations or interactive narratives. Finally, in the Behavior Observation stage, researchers examine emergent social dynamics Zhu et al. (2025)—including cooperation, norm formation, conflict, and human-like heuristic behaviors such as conformity, by observing interactions both among agents and between agents and human participants.
Within this pipeline, Agent Specification is crucial for defining and shaping agent behavior, representative examples include: persona-based specification Rath et al. (2025); Park et al. (2023); Zhou et al. (2024) and role-based specification Li et al. (2023a); Zheng et al. (2023); Chen et al. (2024b); Jia et al. (2024). Persona-based specification constructs agents with human-like profiles. Core elements typically include demographic descriptors (e.g., age, occupation, family structure), personality traits (e.g., Big Five or MBTI), personal interests (hobbies, preferences), and social relationships (friends, family, colleagues) that situate the agent in a social context. In practice, persona detail can range from a single facet (such as demographics) to comprehensive architectures—such as Generative Agents Park et al. (2023)—that integrate multiple attributes with memory to produce situated behavior. Role-based specification defines agents primarily by their professional or task-oriented identities, rather than rich personal backgrounds Chen et al. (2024b). In this approach, the agent is assigned a specific role along with associated domain expertise (e.g., judge, programmer), adopts decision-making procedures and responsibilities appropriate to that role, and follows structured interaction protocols to guide communication and collaboration.
@@ B.1 Example of Persona-Based Specification
”Common” agent specification  Park et al. (2023):
@@ B.2 Example of Role-Based Specification
”Teacher” agent specification Chen et al. (2024b):
@@ C.1 Bandwagon Effect Testbed
Bandwagon effect refers to the psychological tendency for individuals to adopt behaviors, beliefs, or trends simply because they observe others doing so, often motivated by a desire to fit in or belong Asch (1951); Goldstein et al. (2008). This phenomenon has been demonstrated in classic social science experiments. For instance, in the Asch’s Line experiment, participants asked to judge the length of lines were likely to conform to a group’s clearly incorrect answers Asch (1951). In addition, in the Hotel Towel study, hotel guests were more likely to reuse their towels when informed that the majority of previous guests had done so  Goldstein et al. (2008).
@@ 2. Hotel Towel Paradigm: Testing Bandwagon Effect via Majority Behavior
To assess the influence of majority behavior cues on agent choices, we adapt the Hotel Towel study Goldstein et al. (2008). Each scenario presents a situational context and describes the prevailing behavior of the majority.
@@ C.2 Confirmation Bias Testbed
Confirmation bias refers to the tendency to seek, interpret, and recall information in a manner that confirms one’s pre-existing beliefs or hypotheses, while undervaluing or dismissing alternative possibilities Wason (1960); Lord et al. (1979). This phenomenon has been robustly demonstrated in classic social science experiments. For example, in the Wason Selection task, participants typically chose evidence that would confirm a stated rule, rather than seeking information that could potentially disprove it Wason (1960). In addition, the Biased Information study showed that individuals evaluating a controversial issue were more likely to select information supporting their existing views, while disregarding counter-evidence Lord et al. (1979).
@@ 1. Wason Selection Paradigm: Testing Confirmation Bias via Logical Reasoning
To investigate confirmation bias in logical hypothesis testing, we adapt the Wason Selection task Wason (1960). Each scenario provides a rule and several options, challenging the agent to decide which options to test to verify the rule.
@@ 2. Biased Information Seeking Paradigm: Testing Confirmation Bias via Information Selection
To assess confirmation bias in information-gathering, we adapt the Biased Information Seeking study Lord et al. (1979). Each scenario presents a controversial topic and articles with varying alignments, prompting the agent to select which articles to read.
@@ C.3 Framing Effect Testbed
Framing effect refers to the phenomenon where individuals respond differently to identical information depending on whether it is presented as a potential gain or a potential loss Tversky & Kahneman (1981); Kahneman & Tversky (2013). This cognitive bias has been robustly demonstrated in classic experiments. For example, in the Asian Disease study, participants tended to make more risk-averse decisions when outcomes were positively framed (e.g., “200 will be saved”), but exhibited greater risk-seeking behavior when the same outcomes were presented as losses (e.g., “400 will die”) Tversky & Kahneman (1981). Similarly, in the Investment/Insurance study, individuals were more likely to purchase insurance when risks were described as potential losses rather than as missed gains, highlighting how framing can significantly influence decision-making Kahneman & Tversky (2013).
@@ 1. Asian Disease Paradigm: Testing Framing Effect via Risk Decision-Making
To investigate the impact of gain/loss framing on risk preferences, we adapt the Asian Disease study Tversky & Kahneman (1981). Each prompt presents a hypothetical problem and two alternative programs, with outcomes identically described in either a positive (gain) or negative (loss) frame.
@@ 2. Investment/Insurance Paradigm: Testing Framing Effect via Everyday Decision Framing
To assess the influence of positive and negative framing in everyday decision-making, we adapt the Investment/Insurance study Kahneman & Tversky (2013). Each scenario describes a decision context and presents two options, differing only in the framing of equivalent outcomes.

### v2

@@ 9 Limitation
1. Usability: We choose to evaluate CoBRA using technical benchmarks and demonstrations, standard methods for assessing HCI toolkits (Ledo et al., 2018). This paper focuses on exploring the technical feasibility and foundational understanding of the problem. We plan deployment studies with sociologists as the next step to assess CoBRA’s usability (Landay, 2009).
@@ A.1 Cognitive Bias in LLM-As a Feature, Not a Bug
Traditionally, cognitive bias is defined as systematic deviations from rational judgment that lead to suboptimal decision-making outcomes (Unknown, 2015; Baron, 2007). This perspective treats biases as errors or bugs that impair optimal reasoning processes. Recent studies have documented extensive cognitive biases in LLMs, including confirmation bias, anchoring bias, and framing effects (Koo et al., 2023; Boonprakong et al., 2025; Sumita et al., 2025). Research demonstrates that LLMs exhibit amplified cognitive biases in moral decision-making contexts (Cheung et al., 2025), leading to a systematic assessment and mitigation of these biases (Echterhoff et al., 2024).
However, cognitive bias can also be viewed as an adaptive feature that evolved to facilitate social navigation and cooperation (Haselton et al., 2015; Cosmides & Tooby, 1994). From this evolutionary perspective, biases serve important prosocial functions, enabling sophisticated social intelligence rather than representing flaws (Liu et al., 2025b). However, existing research lacks systematic mechanisms to harness these biases as controllable features for enhancing AI social capabilities.
@@ A.2 LLM Behavioral Control Methods
Prompt engineering has served as the foundational approach to behavioral control, leveraging natural language instructions to guide model behavior using structured prompting strategies (Rath et al., 2025; Zhuo et al., 2024). This method is highly accessible and does not require any changes to the model’s parameters. However, designing effective prompts has many pitfalls (Zamfirescu-Pereira et al., 2023), and fundamental limitations persist: LLMs exhibit extreme sensitivity to prompt variations, with large accuracy differences from minor changes (Sclar et al., 2024).
Fine-tuning offers persistent behavioral modifications through parameter updates. Techniques such as Reinforcement Learning from Human Feedback (RLHF) and Direct Preference Optimization (DPO) have become dominant for model alignment (Ouyang et al., 2022; Rafailov et al., 2023). Task vectors enable compositional behavioral modifications through arithmetic operations on weight differences (Ilharco et al., 2023). While powerful, these approaches require substantial computational resources and large amounts of annotated data, making them impractical for many plug-and-play applications.
Representation engineering provides a method for behavioral control through direct manipulation of a model’s internal representations during inference (Zou et al., 2023). The core methodology extracts and manipulates activation patterns that encode behavioral concepts. Activation steering methods compute steering vectors from contrasting behavioral examples (Rimsky et al., 2024), which can be applied during inference to modulate behavior without parameter updates (Todd et al., 2024). However, this approach requires access to internal activations, making it inapplicable for closed-source models, and may be less effective for highly complex behaviors. Recent applications include personality control (Chen et al., 2025), community-specific preference modeling (Li et al., 2025), harmlessness control (Arditi et al., 2024), and multi-objective alignment (Shi et al., 2024). These methods operate with minimal inference overhead while maintaining base model capabilities.
@@ A.3 LLM-based Social Simulation & Agent Specification
Existing work on LLM-based agent social simulations typically instantiates LLMs as autonomous agents by assigning them scenario-specific prompts, roles, and objectives (Park et al., 2023; Chen et al., 2024b). Methodologically, these studies typically follow a pipeline consisting of three core stages: 1. Agent Specification, 2. Social Environment Set-up, and 3. Behavior Observation. During Agent Specification, researchers harness the expressiveness and flexibility of natural language to define agent identity, knowledge, and behavioral tendencies, enabling the creation of contextualized personas. In the Social Environment Set-up stage, agents are embedded in experimental scenarios inspired by social science literature (Liu et al., 2025b), or in constructed social environments with specific rules and resources (Zhou et al., 2024; Park et al., 2023; Chen et al., 2024b), such as sandbox simulations or interactive narratives. Finally, in the Behavior Observation stage, researchers examine emergent social dynamics (Zhu et al., 2025)—including cooperation, norm formation, conflict, and human-like heuristic behaviors such as conformity, by observing interactions both among agents and between agents and human participants.
Within this pipeline, Agent Specification is crucial for defining and shaping agent behavior, representative examples include: persona-based specification (Rath et al., 2025; Park et al., 2023; Zhou et al., 2024) and role-based specification (Li et al., 2023a; Zheng et al., 2023; Chen et al., 2024b; Jia et al., 2024). Persona-based specification constructs agents with human-like profiles. Core elements typically include demographic descriptors (e.g., age, occupation, family structure), personality traits (e.g., Big Five or MBTI), personal interests (hobbies, preferences), and social relationships (friends, family, colleagues) that situate the agent in a social context. In practice, persona detail can range from a single facet (such as demographics) to comprehensive architectures—such as Generative Agents (Park et al., 2023)—that integrate multiple attributes with memory to produce situated behavior. Role-based specification defines agents primarily by their professional or task-oriented identities, rather than rich personal backgrounds (Chen et al., 2024b). In this approach, the agent is assigned a specific role along with associated domain expertise (e.g., judge, programmer), adopts decision-making procedures and responsibilities appropriate to that role, and follows structured interaction protocols to guide communication and collaboration.
@@ B.1 Example of Persona-Based Specification
”Common” agent specification  (Park et al., 2023):
@@ B.2 Example of Role-Based Specification
”Teacher” agent specification (Chen et al., 2024b):
@@ C.1 Bandwagon Effect Testbed
Bandwagon effect refers to the psychological tendency for individuals to adopt behaviors, beliefs, or trends simply because they observe others doing so, often motivated by a desire to fit in or belong (Asch, 1951; Goldstein et al., 2008). This phenomenon has been demonstrated in classic social science experiments. For instance, in the Asch’s Line experiment, participants asked to judge the length of lines were likely to conform to a group’s clearly incorrect answers (Asch, 1951). In addition, in the Hotel Towel study, hotel guests were more likely to reuse their towels when informed that the majority of previous guests had done so  (Goldstein et al., 2008).
@@ 2. Hotel Towel Paradigm: Testing Bandwagon Effect via Majority Behavior
To assess the influence of majority behavior cues on agent choices, we adapt the Hotel Towel study (Goldstein et al., 2008). Each scenario presents a situational context and describes the prevailing behavior of the majority.
@@ C.2 Confirmation Bias Testbed
Confirmation bias refers to the tendency to seek, interpret, and recall information in a manner that confirms one’s pre-existing beliefs or hypotheses, while undervaluing or dismissing alternative possibilities (Wason, 1960; Lord et al., 1979). This phenomenon has been robustly demonstrated in classic social science experiments. For example, in the Wason Selection task, participants typically chose evidence that would confirm a stated rule, rather than seeking information that could potentially disprove it (Wason, 1960). In addition, the Biased Information study showed that individuals evaluating a controversial issue were more likely to select information supporting their existing views, while disregarding counter-evidence (Lord et al., 1979).
@@ 1. Wason Selection Paradigm: Testing Confirmation Bias via Logical Reasoning
To investigate confirmation bias in logical hypothesis testing, we adapt the Wason Selection task (Wason, 1960). Each scenario provides a rule and several options, challenging the agent to decide which options to test to verify the rule.
@@ 2. Biased Information Seeking Paradigm: Testing Confirmation Bias via Information Selection
To assess confirmation bias in information-gathering, we adapt the Biased Information Seeking study (Lord et al., 1979). Each scenario presents a controversial topic and articles with varying alignments, prompting the agent to select which articles to read.
@@ C.3 Framing Effect Testbed
Framing effect refers to the phenomenon where individuals respond differently to identical information depending on whether it is presented as a potential gain or a potential loss (Tversky & Kahneman, 1981; Kahneman & Tversky, 2013). This cognitive bias has been robustly demonstrated in classic experiments. For example, in the Asian Disease study, participants tended to make more risk-averse decisions when outcomes were positively framed (e.g., “200 will be saved”), but exhibited greater risk-seeking behavior when the same outcomes were presented as losses (e.g., “400 will die”) (Tversky & Kahneman, 1981). Similarly, in the Investment/Insurance study, individuals were more likely to purchase insurance when risks were described as potential losses rather than as missed gains, highlighting how framing can significantly influence decision-making (Kahneman & Tversky, 2013).
@@ 1. Asian Disease Paradigm: Testing Framing Effect via Risk Decision-Making
To investigate the impact of gain/loss framing on risk preferences, we adapt the Asian Disease study (Tversky & Kahneman, 1981). Each prompt presents a hypothetical problem and two alternative programs, with outcomes identically described in either a positive (gain) or negative (loss) frame.
@@ 2. Investment/Insurance Paradigm: Testing Framing Effect via Everyday Decision Framing
To assess the influence of positive and negative framing in everyday decision-making, we adapt the Investment/Insurance study (Kahneman & Tversky, 2013). Each scenario describes a decision context and presents two options, differing only in the framing of equivalent outcomes.